OpenStudy (anonymous):

radium-226 has a half life of about 400 years . after 600 years how much radium would be left if you started with 1000 mg?

OpenStudy (anonymous):

find an exponential decay equation for the amount left

OpenStudy (anonymous):

do you know how to do this?

OpenStudy (anonymous):

no

OpenStudy (wolf1728):

Ending Amount = Beginning Amount / 2^n where 'n' is equal to elapsed time / half-life

OpenStudy (wolf1728):

The half life of radium 226 is 1,601 years https://en.wikipedia.org/wiki/Radium

OpenStudy (wolf1728):

1,000 mg after 600 years ? Ending Amount = Beginning Amount / 2^n where 'n' is equal to elapsed time / half-life 'n' = 600/1,601 Ending Amount = 1,000 mg / 2^(0.374765771392879) Ending Amount = 1,000 mg / 1.29662902348742 Ending Amount = 771.23 mg Assuming you had to determine this for substance 'x' which has a half-life of 400 years: Ending Amount = Beginning Amount / 2^n 'n' = 600/400 = 1.5 Ending Amount = 1,000 mg / 2^1.5 Ending Amount = 1,000 mg / 2.82843 Ending Amount = 353.55339 mg